OpenStudy (anonymous):

there are 5 blue pencils and 3 gold pencils in a container. if 2 pencils are drawn at random without replacement, what is the probability that both pencils drawn are gold? express your answer as a fraction in simplest form.

OpenStudy (anonymous):

\[\frac{3}{8}\times \frac{2}{7}\]

OpenStudy (anonymous):

without replacement tells you the draws are independent that allows you to use P(A and B) = P(A)P(B) where A is the probability that the first draw is a gold pencil and B the prob. that the second draw is a gold pencil

OpenStudy (anonymous):

actually without replacement tells you the draws are DEPENDENT

OpenStudy (anonymous):

so it is \(P(A\cap B)=P(A|B)P(B)\)